OpenStudy (elisaneedshelp):

What is the simplest form of ? Expression is provided below Pleasee Medal and Fan !

OpenStudy (elisaneedshelp):

\[8\sqrt[3]{5}-\sqrt[3]{40}-2\sqrt[3]{135}\]

OpenStudy (nincompoop):

do you know the rules for this kind of problem ?

OpenStudy (elisaneedshelp):

please

OpenStudy (elise_a18):

WolframAlpha

OpenStudy (dan815):